# Occupied Processes: Going with the Flow

### v4

@@ 7.1 Optimal Stopping
$v(\mathfrak{o},x)=\sup_{\tau\in\vartheta(\operatorname{\mathcal{T}}_{t})}\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{\mathfrak{o},x}[\varphi(\operatorname{\mathcal{O}}_{\tau},X_{\tau})],\quad\operatorname{\mathcal{T}}_{t}=\operatorname{\mathcal{T}}\cap\ [t,T],\quad t=\mathfrak{o}(\operatorname{\mathbb{R}}).$ (90)
@@ 7.2 Spot Local Time and Early Exercise Premium
$C_{t}=\int_{\operatorname{\mathbb{R}}}L_{t}^{y}\ \phi_{T-t}(y-X_{t})dy+\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[L_{T-t}^{0}],\quad\phi_{s}(x)=\frac{e^{-\frac{x^{2}}{2s}}}{\sqrt{2\pi s}}.$ (95)
@@ 7.3 Numerical Resolution in the case $\operatorname{\mathcal{T}}=[0,T]$
$\tau_{\iota}=\inf\{t\in[\iota,T]\ \mathrel{\mathop{\ordinarycolon}}\ X_{t}=X^{*}_{\iota}\}\wedge T,\quad X^{*}_{\iota}=\text{argmax}_{x}L_{\iota}^{x}.$
@@
$\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[L_{\tau_{\iota}}^{X_{\tau_{\iota}}}]=p_{\iota}\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sup_{x}L_{\iota}^{x}\ |\ \sigma^{*}\geq\iota]+(1-p_{\iota})\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[L_{T}^{X_{T}}\ |\ \sigma^{*}<\iota],\quad p_{\iota}=\operatorname{\mathbb{Q}}(\sigma^{*}\geq\iota).$
@@
$\operatorname{\mathcal{O}}_{t_{n}}=(L_{t_{n}}^{x_{-n}},\ldots,L_{t_{n}}^{x_{n}}),\quad L_{t_{n}}^{x_{m}}=\frac{\operatorname{\mathcal{O}}_{t_{n}}(B_{\varepsilon}(x_{m}))}{2\varepsilon}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2\varepsilon}\sum_{i=1}^{n}\mathds{1}_{\{X_{t_{i}}\ =\ x_{m}\}}\delta t.$ (97)
@@ 7.4 Shrinking the Corridor
$\left\lVert\varphi^{\varepsilon}(\operatorname{\mathcal{O}}_{\tau},X_{\tau})-L_{\tau}^{X_{\tau}}\right\rVert_{L^{2}(\operatorname{\mathbb{Q}})}\leq\sqrt{C\varepsilon}\qquad\forall\tau\in\vartheta(\operatorname{\mathcal{T}}).$ (99)
@@ Appendix A Chronology Invariance
$\psi(s)=t_{\sigma(n)-\xi_{n}^{+}}\ +\ \xi_{n}(s-t_{n-1}),\quad t\in I_{n}\mathrel{\mathop{\ordinarycolon}}=[t_{n-1},t_{n})\quad(I_{N}=[t_{N-1},T]),$ (100)
@@
$\omega_{t}^{\mathfrak{o}}=\sup\{x\in\operatorname{\mathbb{R}}\ \mathrel{\mathop{\ordinarycolon}}\ \mathfrak{o}((-\infty,x])\leq t\},\quad t\leq T.$ (101)
@@ Appendix B Proofs
$Y_{t}-Y_{s}$ $=\sum_{t_{n}\in\Pi^{N}}(Y_{t_{n+1}}-Y_{t_{n}})\quad\qquad(t_{N+1}\mathrel{\mathop{\ordinarycolon}}=t)$
@@
$Z^{N}_{u}=\sum_{t_{n}\in\Pi^{N}}\delta_{\mathfrak{o}}f(\operatorname{\mathcal{O}}_{t_{n}}+\eta_{n}\Delta_{n},X_{t_{n+1}})(X_{u})\ \mathds{1}_{[t_{n},t_{n}+1)}(u),\qquad[t_{N},t_{N+1})=\{t\}.$
@@ B.2 Proposition 4
$\operatorname{\mathcal{O}}_{T}(\omega+h\mathds{1}_{[t,T]})=\operatorname{\mathcal{O}}_{t}(\omega)+\int_{t}^{T}\delta_{\omega_{s}+h}ds=\operatorname{\mathcal{O}}_{t}(\omega)+\operatorname{\mathcal{O}}_{t,T}^{h}(\omega),\quad\forall A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}),$
@@ B.5 Proposition 6
$v^{\varepsilon}(T)=\frac{2}{\varepsilon}\int_{0}^{\varepsilon}c(T,x)dx=v(T)-\frac{\varepsilon}{2}+\frac{\varepsilon^{2}/3}{\sqrt{2\pi T}}(1+r_{\varepsilon}),\quad\lim_{\varepsilon\downarrow 0}r_{\varepsilon}=0.$
@@ B.6 Proposition 7
$\left\lVert\sup_{(t,x)\in[0,T]\times\operatorname{\mathbb{R}}}|L_{t}^{x+z}-L_{t}^{x}|\right\rVert_{L^{2}(\operatorname{\mathbb{Q}})}^{2}\leq C|z|,\quad|z|\leq 1.$
@@ C.1 Least Square Monte Carlo for Spot Local Time
    Space-time grid $\begin{cases}t_{n}\ =n\delta t,\;\ &n=0,\ldots,N,\quad\delta t=\frac{T}{N},\\ x_{m}=m\sqrt{3\delta t},\;\ &m=-N,\ldots,N\end{cases}$
@@
    $X_{t_{n+1}}=X_{t_{n}}+Z_{n+1}\sqrt{\delta t},\quad X_{0}=0,\quad Z_{1},\ldots,Z_{N}\overset{i.i.d.}{\sim}\nu=\frac{1}{6}\left(\delta_{-\sqrt{3}}+4\delta_{0}+\delta_{\sqrt{3}}\right)$

### v5

@@ 7.1 Optimal Stopping
$v(\mathfrak{o},x)=\sup_{\tau\in\vartheta(\operatorname{\mathcal{T}}_{t})}\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{\mathfrak{o},x}[\varphi(\operatorname{\mathcal{O}}_{\tau},X_{\tau})],\qquad\operatorname{\mathcal{T}}_{t}=\operatorname{\mathcal{T}}\cap\ [t,T],\qquad t=\mathfrak{o}(\operatorname{\mathbb{R}}).$ (91)
@@ 7.2 Spot Local Time and Early Exercise Premium
$C_{t}=\int_{\operatorname{\mathbb{R}}}L_{t}^{y}\ \phi_{T-t}(y-X_{t})dy+\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[L_{T-t}^{0}],\qquad\phi_{s}(x)=\frac{e^{-\frac{x^{2}}{2s}}}{\sqrt{2\pi s}}.$ (96)
@@ 7.3 Numerical Resolution in the case $\operatorname{\mathcal{T}}=[0,T]$
$\tau_{\iota}=\inf\{t\in[\iota,T]\ \mathrel{\mathop{\ordinarycolon}}\ X_{t}=X^{*}_{\iota}\}\wedge T,\qquad X^{*}_{\iota}=\text{argmax}_{x}L_{\iota}^{x}.$
@@
$\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[L_{\tau_{\iota}}^{X_{\tau_{\iota}}}]=p_{\iota}\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sup_{x}L_{\iota}^{x}\ |\ \sigma^{*}\geq\iota]+(1-p_{\iota})\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[L_{T}^{X_{T}}\ |\ \sigma^{*}<\iota],\qquad p_{\iota}=\operatorname{\mathbb{Q}}(\sigma^{*}\geq\iota).$
@@
$\operatorname{\mathcal{O}}_{t_{n}}=(L_{t_{n}}^{x_{-n}},\ldots,L_{t_{n}}^{x_{n}}),\qquad L_{t_{n}}^{x_{m}}=\frac{\operatorname{\mathcal{O}}_{t_{n}}(B_{\varepsilon}(x_{m}))}{2\varepsilon}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2\varepsilon}\sum_{i=1}^{n}\mathds{1}_{\{X_{t_{i}}\ =\ x_{m}\}}\delta t.$ (98)
@@ 7.4 Shrinking the Corridor
$\left\lVert\varphi^{\varepsilon}(\operatorname{\mathcal{O}}_{\tau},X_{\tau})-L_{\tau}^{X_{\tau}}\right\rVert_{L^{2}(\operatorname{\mathbb{Q}})}\leq\sqrt{C\varepsilon}\hskip 17.00024pt\forall\tau\in\vartheta(\operatorname{\mathcal{T}}).$ (100)
@@ Appendix A Chronology Invariance
$\psi(s)=t_{\sigma(n)-\xi_{n}^{+}}\ +\ \xi_{n}(s-t_{n-1}),\qquad t\in I_{n}\mathrel{\mathop{\ordinarycolon}}=[t_{n-1},t_{n})\qquad(I_{N}=[t_{N-1},T]),$ (101)
@@
$\omega_{t}^{\mathfrak{o}}=\sup\{x\in\operatorname{\mathbb{R}}\ \mathrel{\mathop{\ordinarycolon}}\ \mathfrak{o}((-\infty,x])\leq t\},\qquad t\leq T.$ (102)
@@ Appendix B Proofs
$Y_{t}-Y_{s}$ $=\sum_{t_{n}\in\Pi^{N}}(Y_{t_{n+1}}-Y_{t_{n}})\qquad\hskip 17.00024pt(t_{N+1}\mathrel{\mathop{\ordinarycolon}}=t)$
@@
$Z^{N}_{u}=\sum_{t_{n}\in\Pi^{N}}\delta_{\mathfrak{o}}f(\operatorname{\mathcal{O}}_{t_{n}}+\eta_{n}\Delta_{n},X_{t_{n+1}})(X_{u})\ \mathds{1}_{[t_{n},t_{n}+1)}(u),\hskip 17.00024pt[t_{N},t_{N+1})=\{t\}.$
@@ B.2 Proposition 4
$\operatorname{\mathcal{O}}_{T}(\omega+h\mathds{1}_{[t,T]})=\operatorname{\mathcal{O}}_{t}(\omega)+\int_{t}^{T}\delta_{\omega_{s}+h}ds=\operatorname{\mathcal{O}}_{t}(\omega)+\operatorname{\mathcal{O}}_{t,T}^{h}(\omega),\qquad\forall A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}),$
@@ B.5 Proposition 7
$v^{\varepsilon}(T)=\frac{2}{\varepsilon}\int_{0}^{\varepsilon}c(T,x)dx=v(T)-\frac{\varepsilon}{2}+\frac{\varepsilon^{2}/3}{\sqrt{2\pi T}}(1+r_{\varepsilon}),\qquad\lim_{\varepsilon\downarrow 0}r_{\varepsilon}=0.$
@@ B.6 Proposition 8
$\left\lVert\sup_{(t,x)\in[0,T]\times\operatorname{\mathbb{R}}}|L_{t}^{x+z}-L_{t}^{x}|\right\rVert_{L^{2}(\operatorname{\mathbb{Q}})}^{2}\leq C|z|,\qquad|z|\leq 1.$
@@ C.1 Least Square Monte Carlo for Spot Local Time
    Space-time grid $\begin{cases}t_{n}\ =n\delta t,\;\ &n=0,\ldots,N,\qquad\delta t=\frac{T}{N},\\ x_{m}=m\sqrt{3\delta t},\;\ &m=-N,\ldots,N\end{cases}$
@@
    $X_{t_{n+1}}=X_{t_{n}}+Z_{n+1}\sqrt{\delta t},\qquad X_{0}=0,\qquad Z_{1},\ldots,Z_{N}\overset{i.i.d.}{\sim}\nu=\frac{1}{6}\left(\delta_{-\sqrt{3}}+4\delta_{0}+\delta_{\sqrt{3}}\right)$